OpenStudy (anonymous):

if f(x)=8sqrtx, find f'(x) and f"(x) if f(x)=3rd root of sqrtx^4, find f'(x) and f"(x) help :(

zepdrix (zepdrix):

\[\Large\rm f(x)=8\sqrt x\]\[\Large\rm f(x)=8x^{1/2}\]Ok with that step?

zepdrix (zepdrix):

Rewriting a root as a rational exponent allows you to easily apply your power rule.

OpenStudy (anonymous):

zepdrix you got this one?

zepdrix (zepdrix):

mhm

OpenStudy (anonymous):

yes I understand that step

zepdrix (zepdrix):

So what do you get for your first derivative? :)\[\Large\rm f'(x)=8\cdot\frac{1}{2}x^{\frac{1}{2}-1}\]Something like that, yah? Understand how to simplify the exponent?

OpenStudy (anonymous):

is it -4x^1/2 ?

OpenStudy (anonymous):

oops i mean to the -1/2

OpenStudy (anonymous):

okay so f'(x)=4x^-1/2 & f"(x)=-2x^-3/2 ?

zepdrix (zepdrix):

Mmm yes good! :)

zepdrix (zepdrix):

Next one will follow the same process, power rule and such. We just need to get it into the proper rationalize exponent form.\[\Large\rm f(x)=\sqrt[3]{x^4}\]

zepdrix (zepdrix):

The degree of the root becomes the denominator. In our last question, we had a `square root`, which is a `2nd root`. That's why we had a 2 on the bottom of our fraction.

OpenStudy (anonymous):

okay so f'(x) will be 4/3x^1/3?

zepdrix (zepdrix):

good c:

OpenStudy (anonymous):

thanks :) & f"(x)=4/9x^-2/3

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

lol thank you :) I just needed some reassurance. I appreciate your help :)